Solve. Please show all of your work.
A triangular lot has a perimeter of 1900 feet. One side is 200 feet longer than the shortest side, while the third side is 500 feet longer than the shortest side. Find the lengths of all three sides

Answer by Theo(13342)

let the sides be x,y,z.
let side x be the shortest side.
y = 200 + x
z = 500 + x
the perimeter = 1900 feet.
x + y + z = 1900
substitute 200 + x for y to get:
x + 200 + x = 1900
substitute 500 + x for z to get:
x + 200 + x + 500 + x = 1900
combine like terms to get:
3x + 700 = 1900
subtract 700 from both sides of the equation to get:
3x = 1200
divide both sides by 3 to get:
x = 400
this makes:
x = 400
y = 600
z = 900
x + y + z = 1900
the answer is:
x = 400
y = 600
z = 900
x = shortest side
y = 200 + x
z = 500 + x